# Toward Relative Positional Encoding in Spiking Transformers

### v1

@@ Abstract
Spiking neural networks (SNNs) are bio-inspired networks that model how neurons in the brain communicate through discrete spikes, which have great potential in various tasks due to their energy efficiency and temporal processing capabilities. SNNs with self-attention mechanisms (Spiking Transformers) have recently shown great advancements in various tasks such as sequential modeling and image classifications. However, integrating positional information, which is essential for capturing sequential relationships in data, remains a challenge in Spiking Transformers. In this paper, we introduce an approximate method for relative positional encoding (RPE) in Spiking Transformers, leveraging Gray Code as the foundation for our approach. We provide comprehensive proof of the method’s effectiveness in partially capturing relative positional information for sequential tasks. Additionally, we extend our RPE approach by adapting it to a two-dimensional form suitable for image patch processing. We evaluate the proposed RPE methods on several tasks, including time series forecasting, text classification, and patch-based image classification. Our experimental results demonstrate that the incorporation of RPE significantly enhances performance by effectively capturing relative positional information.
Spiking Neural Networks, Relative Positional Encoding
@@ 1 Introduction
Spiking Neural Networks (SNNs) (Maass, 1997) are a class of bio-inspired models designed to emulate the communication process of biological neurons, which transmit information through discrete spikes. In contrast to artificial neural networks (ANNs) that operate on continuous values, SNNs process information in the form of spikes occurring at precise moments in time. This discrete event-driven processing enables SNNs to be highly energy-efficient. The temporal characteristics of spikes make SNNs particularly well-suited for tasks involving sequential data or dynamic environments, such as sensory processing (Li et al., 2017a; Fang et al., 2021), patch-based image classification (Zhou et al., 2023b; Yao et al., 2023), time-series forecasting (Lv et al., 2024b; Park et al., 2024), and natural language processing (Zhu et al., 2023; Lv et al., 2023b; Xing et al., 2024).
In ANNs, relative positional encoding (RPE) is often used to encode the positions of input elements (Su et al., 2024), enabling the model to capture the relative relationships between positions within a sequence. While RPE has been shown to enhance performance in language modeling (Su et al., 2024) and image processing (Liu et al., 2021), its direct application to SNNs remains limited due to the unique characteristics of spike-based computation. Current approaches to realize positional encoding (PE) in SNNs either fail to maintain the uniqueness of spike representations for each position (Zhou et al., 2023b; Yao et al., 2024a), or ignore the incorporation of relative positional information (Lv et al., 2024a).
In this paper, we first identify a limitation in the current spiking self-attention (SSA) mechanism (Zhou et al., 2023b), which is suboptimal for measuring the distance (or similarity) between tokens in the query and key matrices. Specifically, the dot product operation fails to effectively differentiate distances when both the query and key are spike matrices. To address this limitation, we propose a new spiking self-attention mechanism based on the Exclusive-NOR (XNOR) operation, which enhances the measurement of query-key similarity by leveraging the Hamming distance (see Section 3.4).
Secondly, we introduce a novel approximate relative positional encoding method for SNNs, based on Gray Code (Gray, 1953) as the foundational encoding scheme. The consistent Hamming distance between Gray Code representations of two position indices differing by $2^{n}$ allows for serving as a spiking RPE method and the integration with the XNOR-based SSA mechanism. The binary encoding nature of Gray Code is well-suited to represent relative positions while preserving the spiking dynamics of SNNs, which is crucial for effectively processing sequential data. Furthermore, we provide a theoretical proof in Appendix A demonstrating the efficacy of Gray Code in partially capturing relative positional information for sequential tasks.
Moreover, we extend our RPE method to the logarithmic form that can be directly integrated into the attention map of XNOR-based spiking self-attention, offering greater efficiency compared to the previous form while maintaining comparable performance. For image patches, we also design a two-dimensional version of our method to capture the relative positional information of both height and width.
To evaluate the effectiveness of our RPE approaches, we apply them to three diverse tasks: time series forecasting, text classification, and patch-based image classification. Our experiments demonstrate that the proposed RPE method enhances the performance of SNNs across all three tasks, allowing the original SNNs to capture and leverage relative relationships between elements in the input sequential data. Moreover, we further conduct an ablation study and positional encoding analysis to validate the inner properties of our proposed RPE method. Based on the analysis, we offer several valuable insights into designing effective positional encoding schemes for SNNs.
We believe that this work contributes a new method for integrating relative positional encoding into SNNs, expanding their potential for a wide range of applications in machine learning and computational neuroscience. To summarize, our contributions are as follows:
  • A Novel RPE Method for SNNs. Our research represents one of the pioneering efforts to propose a solution to relative positional encoding for spiking neural networks, albeit as an approximation method.
@@
  • Consistent Performance Improvement. Our proposed method and its extended variants significantly and consistently improve the performance of SNNs across various sequential tasks, including time-series forecasting and text classification.
  • Theoretical Validation and Insightful Analysis. In addition to experimental evaluations, we provide thorough theoretical proof to validate the effectiveness of our RPE method. Furthermore, we offer an analysis of the underlying properties of RPE, providing valuable insights for future designs of RPE in SNNs.
@@ 2 Related Work
In sequential tasks, positional encoding (PE) plays a crucial role in capturing the order of tokens. Traditional absolute positional encoding assigns a fixed position to each token, whereas relative positional encoding (RPE) emphasizes the distance between tokens. RPE allows the model to generalize across different sequence lengths and better capture relationships between tokens.
Despite its significance, there has been limited research focusing on the role of positional encoding approaches in spiking neural networks. Notably, both Spikformer (Zhou et al., 2023b) and Spike-driven Transformer (Yao et al., 2023, 2024a) incorporate a combination of convolutional layers, batch normalization, and spiking neuron layers to generate “learnable relative positional encodings”. However, we argue that this approach functions more similarly to a spike-element-wise residual connection (Fang et al., 2021) than to a conventional positional encoding module. A robust PE module should offer unique representations for each position, but the spike-position matrices generated by these methods may lead to identical spike representations for different positions.
Lv et al. (2024a) introduced a spiking positional encoding method for SNNs, inspired by central pattern generators (Marder & Bucher, 2001). This method generates periodic positional encoding, ensuring a unique binary spike matrix for each position. However, their approach is based on absolute positional encoding and, thus, does not capture the time-translational invariance property in many sequential modeling problems, which, however, is an important advantage of relative positional encoding methods.
@@ 3 Preliminary
Figure 1: Preliminary Figures. (a) Spike dynamics of spiking (LIF) neurons. (b) Illustration of spiking self-attention in Spikformer (Zhou et al., 2023b) and QKFormer (Zhou et al., 2024). (c) An example of Hamming Distance. (d) Calculation process of Gray Code.
3.1 Spiking Neural Networks
The fundamental building block of SNNs is the spiking neuron, such as the leaky integrate-and-fire (LIF) model (Maass, 1997), which is governed by the input current $I(t)$, influencing the membrane potential $U(t)$ and the spike output $S(t)$ at each time step $t$. The dynamics of the LIF neuron are captured by the following system of equations:
$U(t)=H(t-\Delta t)+I(t),\quad I(t)=f(\mathbf{x};\mathbf{\theta}),$ (1)
$H(t)=V_{\rm reset}S(t)+(1-S(t))\beta U(t),$ (2)
$S(t)=\begin{cases}1,&\text{if }U(t)\geq U_{\rm thr},\\ 0,&\text{if }U(t)<U_{\rm thr}.\end{cases}$ (3)
Here, $I(t)$ represents the spatial input to the LIF neuron at time $t$, which is computed by the function $f$ using the input vector $\mathbf{x}$ and the learnable parameters $\mathbf{\theta}$. The discretization constant $\Delta t$ controls the time granularity in the LIF model, and $H(t)$ is the temporal output at time step $t$. The spike $S(t)$ is determined by a Heaviside step function that is activated when the membrane potential $U(t)$ exceeds a threshold $U_{\rm thr}$. Upon reaching the threshold, the neuron generates a spike, and $H(t)$ resets to $V_{\rm reset}$. If $U(t)$ does not exceed the threshold, the membrane potential decays to $H(t)$ at a rate determined by $\beta$.
@@ 3.2 Spiking Self-Attention
Spiking self-attention (SSA) is a spiking version of self-attention (Vaswani, 2017), which was proposed in Spikformer (Zhou et al., 2023b) and QKFormer (Zhou et al., 2024). The vital design is to utilize discrete spikes to approximate the vanilla self-attention mechanism. It can be written as:
$\small\begin{split}\mathbf{Q}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{Q}}\right)\right),\\ \mathbf{K}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{K}}\right)\right),\\ \mathbf{V}&=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W_{V}}\right)\right),\end{split}$ (4)
where $\mathcal{SN}$ is a spike neuron layer described in Equation 3. The input is denoted as $\mathbf{X}\in\{0,1\}^{T\times L\times D}$, where $T$ is the number of time steps. $\operatorname{BN}$ represents batch normalization, and $\tau$ is a scaling factor. The attention map $\mathbf{AttnMap}$ is then computed as the dot product between $\mathbf{Q}$ and $\mathbf{K}^{\operatorname{T}}$:
$\small\mathbf{AttnMap}=\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}},$ (5)
$\small\operatorname{SSA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\mathcal{SN}(\operatorname{BN}((\mathbf{Attn}\cdot\mathbf{V}*\tau)\cdot\mathbf{W})).$ (6)
@@ 3.3 Relative Positional Encoding
In self-attention mechanisms, the attention weights are modified using a relative positional encoding (RPE) $\mathbf{R}_{i,j}$, which depends on the distance between the $i$-th and $j$-th tokens. A typical realization of RPE (Liu et al., 2021; Hao et al., 2024) is formulated as:
$\small\begin{split}\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})&=\mathbf{AttnMap\cdot V}\\ &=\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}}{\sqrt{d_{k}}}+\mathbf{R}_{i,j}\right)\mathbf{V}.\end{split}$ (7)
Here, $\mathbf{R}_{i,j}$ represents the relative distance between tokens $i$ and $j$. Specifically, the relative positional encoding for the $i$-th position in $\mathbf{Q}$ and the $j$-th position in $\mathbf{K}$ should satisfy distance consistency, i.e., the magnitude of $\mathbf{R}_{i,j}$ should depend solely on the relative distance $|i-j|$, enabling the model to distinguish between nearby and distant tokens. This approach improves the model’s ability to handle long-range dependencies and sequence variations.
Figure 2: Overview of Our Method. (a) XNOR-based Spiking Self-Attention (SSA). This panel illustrates the computation flow for the query ($\mathbf{Q}$) and key ($\mathbf{K}$) matrices in a PyTorch-style notation. (b) Gray-PE. Gray code representations exhibit a consistent Hamming distance for two integers differing by $2^{n}$. Gray-PE is implemented by concatenating $G(\bm{l})$ along the $D$ dimension of both $\mathbf{Q}$ and $\mathbf{K}$. (c) Log-PE. A pre-assigned relative distance encoding map $\mathbf{R}_{i,j}\in\mathbb{N}_{0}$ is added to the standard attention map $\mathbf{AttnMap}$. The shading indicates the magnitude of the values, with deeper blue corresponding to larger values. (d) 2D Form of Gray-PE. For image patches, a 2D relative positional encoding is more suitable than the 1D version, as it better captures the spatial relationships.
@@ 3.4 Hamming Distance
The Hamming distance between two binary strings of equal length is the number of bit positions at which the corresponding bits are different. Formally, for two binary strings $A$ and $B$ of length $m$,
$\small d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),$ (8)
where
$\small\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (9)
@@ 4 Method
4.1 Gray Code Definition
Gray Codes (Gray, 1953), also known as reflected binary codes, are binary sequences where consecutive numbers differ by exactly one bit. Gray Code $G(x)$ for a non-negative integer $x$ can be generated using the formula: The standard binary reflected Gray Code $G(x)$ for a non-negative integer $x$ is defined as:
$\small G(x)=x\oplus\left(x\gg 1\right),$ (10)
where $\oplus$ denotes the bitwise XOR operation, and $\gg$ denotes the arithmetic right shift (with zero padding).

### v2

@@ Abstract
Spiking neural networks (SNNs) are bio-inspired networks that mimic how neurons in the brain communicate through discrete spikes, which have great potential in various tasks due to their energy efficiency and temporal processing capabilities. SNNs with self-attention mechanisms (spiking Transformers) have recently shown great advancements in various tasks, and inspired by traditional Transformers, several studies have demonstrated that spiking absolute positional encoding can help capture sequential relationships for input data, enhancing the capabilities of spiking Transformers for tasks such as sequential modeling and image classification. However, how to incorporate relative positional information into SNNs remains a challenge. In this paper, we introduce several strategies to approximate relative positional encoding (RPE) in spiking Transformers while preserving the binary nature of spikes. Firstly, we formally prove that encoding relative distances with Gray Code ensures that the binary representations of positional indices maintain a constant Hamming distance whenever their decimal values differ by a power of two, and we propose Gray-PE based on this property. In addition, we propose another RPE method called Log-PE, which combines the logarithmic form of the relative distance matrix directly into the spiking attention map. Furthermore, we extend our RPE methods to a two-dimensional form, making them suitable for processing image patches. We evaluate our RPE methods on various tasks, including time series forecasting, text classification, and patch-based image classification, and the experimental results demonstrate a satisfying performance gain by incorporating our RPE methods across many architectures. Our results provide fresh perspectives on designing spiking Transformers to advance their sequential modeling capability, thereby expanding their applicability across various domains.
@@ 1 Introduction
Spiking Neural Networks (SNNs) [1] are a class of bio-inspired models designed to emulate the communication process of biological neurons, which transmit information through discrete spikes. In contrast to artificial neural networks (ANNs) that operate on continuous values, SNNs process information in the form of spikes occurring at precise moments in time. The temporal characteristics of spikes make SNNs particularly well-suited for tasks involving sequential data or dynamic environments, such as sensory processing [2, 3], patch-based image classification [4, 5], time-series forecasting [6, 7, 8], and natural language processing [9, 10, 11].
In the vanilla Transformer architecture [12], positional encoding serves as a critical mechanism for modeling sequential dependencies in input data. Beyond absolute positional encoding, relative positional encoding (RPE) [13, 14] has emerged as an effective approach to represent inter-element distances, enabling models to capture relational patterns within sequences dynamically. Although RPE has demonstrated effectiveness in improving language modeling [13] and visual recognition tasks [15], its integration into SNNs remains underexplored. Existing methodologies for implementing positional encoding in spiking Transformers either suffer from ambiguous spike representations across positions [4, 16], or neglect to integrate relative positional relationships entirely [7]. Directly adapting current RPE techniques, such as Attention with Linear Biases (ALiBi) [13] and Rotary Position Embedding (RoPE) [14], to spiking Transformers encounters significant challenges. Specifically, spiking neural architectures exhibit intrinsic difficulty in decoupling relative positional information from their sparse, event-driven representations. This limitation, empirically demonstrated in Section 5.2, underscores the necessity for rethinking RPE integration to align with neuromorphic computing principles, such as temporal sparsity and spike-based communication.
In this paper, we first propose that the Hamming distance [17], which quantifies the number of ones resulting from the XOR operation between two binary strings, serves as an appropriate metric for measuring relative distances when both the query and key matrices are binary. Consequently, we refine the spiking self-attention mechanism [4] by replacing dot-product operations with exclusive-NOR (XNOR) logic operations. Then we present two novel approximation strategies for integrating RPE into spiking Transformers, while strictly preserving the binary activation dynamics inherent to spiking neurons. First, we propose Gray-PE, a method exploiting the properties of Gray Code [18] to binarize relative positional distances. We theoretically prove that encoding relative distances via Gray Code ensures a constant Hamming distance between the binary representations of positional indices whose decimal differences equal $2^{n}$, where $n\geq 0$ (See Theorem 1). This property guarantees that any pair of positions separated by a relative distance of $2^{n}$ in decimal space exhibits invariant Hamming distances in their Gray Code-encoded representations. Such invariance stabilizes positional relationship modeling for power-of-two intervals, addressing a critical limitation in existing spiking neural architectures. Second, we propose Log-PE, a method adapting insights from ALiBi [13] and Rectified RoPE [19]. Log-PE integrates a non-negative logarithmic transformation of the relative distance map directly into the spiking attention map, inducing a decaying sensitivity to positional relationships akin to windowed attention mechanisms. Moreover, we extend the proposed RPE methods to their two-dimensional form, making them suitable for processing image patches.
To systematically evaluate the efficacy of our proposed RPE methods, we benchmark them across three cross-domain tasks: time series forecasting, text classification, and patch-based image classification. We employ three representative spiking Transformer architectures as backbones: Spikformer [4], the Spike-driven Transformer [5], and QKFormer [20]. Experimental results demonstrate consistent performance gains across all tasks when integrating our RPE approaches, affirming that explicit modeling of relative positional relationships addresses a critical limitation in existing spiking Transformer designs. Furthermore, we conduct experiments on ablation study, long sequence modeling, and sensitivity analysis to validate the inner properties of our proposed RPE method.
This work establishes a framework for integrating relative positional encoding (RPE) into spiking Transformers, advancing their applicability in neuro-inspired machine learning paradigms. Our primary contributions are summarized as follows:
  • Two RPE Methods for Spiking Transformers. To our knowledge, this study is among the first to explore RPE adaptations for spiking architectures systematically. While Gray-PE and Log-PE operate as principled approximations constrained by binary spike dynamics, they address a critical gap in positional modeling for neuromorphic computation.
  • Theoretical Foundations and Empirical Analysis. In addition to empirical validation, we provide theoretical guarantees demonstrating that our methods can partially encode relative positional information. Furthermore, we offer necessary analysis on the internal properties of RPE and their robustness facing long sequences.
@@
  • Consistent Performance Gains Across Architectures and Tasks. Our proposed RPE methods consistently improve the performance of spiking Transformers across various sequential tasks, including time-series forecasting and text classification.
@@ 2 Related Work
Positional encoding serves as an indispensable mechanism for preserving the order of input elements in sequential modeling tasks. Traditional absolute positional encoding assigns static, predefined embeddings to individual tokens based on their sequential indices. In contrast, relative positional encoding (RPE) dynamically models the pairwise distances between tokens, enabling the self-attention mechanism to prioritize interactions based on their relative proximity. RPE allows the model to generalize across different sequence lengths and better capture relationships between tokens.
Despite the importance of PE in sequence-aware architectures, its application to SNNs is limited. Existing implementations, such as Spikformer [4] and Spike-driven Transformer [5, 16, 21], incorporate a combination of convolutional layers, batch normalization, and spiking neuron layers to derive learnable positional encodings. However, we argue that this approach functions more similarly to a spike-element-wise residual connection [3] than to a conventional positional encoding module. A principled PE module should offer unique representations for positions, but the spike-position matrices generated by these methods may lead to identical spike representations for different positions.
CPG-PE, proposed by [7], introduces a spiking absolute positional encoding inspired by central pattern generators [22], generating unique periodic binary spike patterns for each position. However, their approach is based on absolute positional encoding and, thus, does not capture the time-translational invariance property in many sequential modeling problems, which, however, is an important advantage of relative positional encoding methods.
@@ 3 Preliminary
Figure 1: Illustration of preliminary knowledge. (a) Spike dynamics of LIF neurons. (b) Illustration of vanilla spiking self-attention in Spikformer [4]. (c) An example of Hamming Distance between two spike trains. (d) The calculation process of the classic Reflected Gray Code.
3.1 Spiking Neurons
We take the leaky integrate-and-fire (LIF) neuron [1] as our building brick of SNNs, which is governed by the input current $I[t]$, influencing the membrane potential $U[t]$ and the spike output $S[t]$ at each time step $t$. The dynamic of the LIF neuron is captured by the following system of equations:
$U[t]$ $=H[t](1-S[t])+U_{\text{reset}}S[t],\quad S[t]=\Theta(H[t]-U_{\text{thr}}),$ (1)
$H[t]$ $=U[t-1]+\frac{1}{\tau}(I[t]-(U[t-1]-U_{\text{reset}})),$ (2)
where $\tau$ is the membrane time constant. The spike $S(t)$ will be triggered when the membrane potential $H(t)$ exceeds a threshold $U_{\text{thr}}$, right after which $U[t]$ will be reset to $U_{\text{reset}}$.
@@ 3.2 Spiking Self-Attention
Spiking self-attention (SSA) is a spiking version of self-attention [12], which was proposed in Spikformer [4]. The vital design is to utilize discrete spikes to approximate the vanilla self-attention mechanism. It can be written as:
$\mathbf{Q},\mathbf{K},\mathbf{V}$ $=\mathcal{SN}\left(\operatorname{BN}\left(\mathbf{X}\cdot\mathbf{W}_{{Q,K,V}}\right)\right)\in\{0,1\}^{T\times L\times D}$ (3)
where $\mathcal{SN}$ is a spike neuron layer described in Equation 1. The input is denoted as $\mathbf{X}\in\{0,1\}^{T\times L\times D}$, where $T$ is the number of time steps. $\operatorname{BN}$ represents batch normalization, and $\sigma$ is a scaling factor. The attention map $\mathbf{AttnMap}$ is then computed as the dot product between $\mathbf{Q}$ and $\mathbf{K}^{\operatorname{T}}$:
$\operatorname{SSA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\mathcal{SN}(\operatorname{BN}((\underbrace{\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}}_{\mathbf{AttnMap}}\cdot\mathbf{V}*\sigma)\cdot\mathbf{W})).$ (4)
@@
Recent studies, including Spike-Driven Transformer (SDT) [5, 16, 21], SpikingResFormer [23], and QKFormer [20], have proposed various modifications to the standard SSA mechanism. For our empirical evaluation, we selectively employ architectures demonstrating compatibility with our proposed relative position encoding methods.
@@ 3.3 Relative Positional Encoding
Relative positional encoding (RPE) in Transformers primarily introduces bias terms into the self-attention mechanism that dynamically encode pairwise token distances. A common implementation of RPE, as demonstrated in prior work [15, 24], is formalized as follows:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\underbrace{\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}}{\sqrt{d_{k}}}+\mathbf{R}_{i,j}\right)}_{\mathbf{AttnMap}}\cdot\mathbf{V}.$ (5)
Here, $\mathbf{R}_{i,j}$ represents the relative positional bias between the $i$-th query and the $j$-th key positions. A critical aspect of RPE is its adherence to distance consistency: the magnitude of $\mathbf{R}_{i,j}$ is determined exclusively by the relative positional offset $|i-j|$, ensuring that the model systematically differentiates between proximal and distant tokens. This property enhances the model’s capacity to capture long-range dependencies and generalize across variations in sequence length and structure.
@@ 3.4 Hamming Distance
The Hamming distance [17] between two binary strings of equal length is the number of bit positions at which the corresponding bits are different. Formally, for two binary strings $A$ and $B$ of length $m$,
$\small d_{H}(A,B)=\sum_{i=1}^{m}\delta(A_{i},B_{i}),\quad\text{where}\quad\delta(A_{i},B_{i})=\begin{cases}1&\text{if }A_{i}\neq B_{i},\\ 0&\text{otherwise}.\end{cases}$ (6)
Hamming distance is suitable for measuring the relative distances when $\mathbf{Q}$ and $\mathbf{K}$ are spike matrices.
3.5 Gray Code
Gray Codes [18], also known as reflected binary codes, are binary numbering systems where adjacent values differ by precisely one bit. For a non-negative integer $x$, the standard binary reflected Gray Code $G(x)$ is defined by the following bitwise operation:
$\small G(x)=x\oplus\left(x\gg 1\right),$ (7)
where $\oplus$ denotes the bitwise XOR operation, and $\gg$ denotes the arithmetic right shift.
@@ 4 Method
Figure 2: Overview of Our Method. (a) XNOR-based spiking self-attention. We illustrate the computation flow for $\mathbf{Q}$ and $\mathbf{K}$ in a PyTorch-style notation. (b) Gray-PE. Position indices differing by $2^{n}$ exhibit a consistent Hamming distance on their Gray code representations. Gray-PE is implemented by concatenating $G(\bm{l})$ along the $D$ dimension on both $\mathbf{Q}$ and $\mathbf{K}$. (c) Log-PE. A pre-assigned relative distance encoding map $\mathbf{R}_{i,j}\in\mathbb{N}_{0}$ is added to the original attention map $\mathbf{AttnMap}$. (d) 2D Form of Gray-PE. A 2D RPE is more suitable than the 1D version for image patches, as it captures the spatial relationships more effectively.
4.1 Design Principles
Relative position encoding (RPE) aims to encode the relative distances between positional indices within a sequence. In many spiking Transformers, such as Spikformer, Spike-Driven Transformer, and QKFormer, both the $\mathbf{Q}$ and $\mathbf{K}$ matrices are binary. Consequently, their relative distances can be computed using the Hamming distance, which corresponds to the number of ones resulting from the XOR operation between $\mathbf{Q}$ and $\mathbf{K}$. To better align with this Hamming distance-based similarity measure, we replace the traditional dot-product spiking self-attention (SSA) mechanism with an XNOR-based SSA. Inspired by RPE strategies in Transformers, we propose two approaches for incorporating relative distance information into spiking attention mechanisms: (1) Gray-PE: Gray-Code-based positional encoding concatenated to $\mathbf{Q}$ and $\mathbf{K}$, and (2) Log-PE: logarithmic positional encoding applied directly to the attention map.